# Clustering in Causal Attention Masking

### v1

@@ Abstract
This work presents a modification of the self-attention dynamics proposed by Geshkovski et al., 2023b to better reflect the practically relevant, causally masked attention used in transformer architectures for generative AI. This modification translates into an interacting particle system that cannot be interpreted as a mean-field gradient flow. Despite this loss of structure, we significantly strengthen the results of Geshkovski et al., 2023b in this context: While previous rigorous results focused on cases where all three matrices (Key, Query, and Value) were scaled identities, we prove asymptotic convergence to a single cluster for arbitrary key-query matrices and a value matrix equal to the identity. Additionally, we establish a connection to the classical Rényi parking problem from combinatorial geometry to make initial theoretical steps towards demonstrating the existence of meta-stable states.
@@ 1   Introduction
Following Sander et al., (2022), recent studies by Geshkovski et al., 2023a and Geshkovski et al., 2023b have proposed a mathematical framework to analyze Transformers as interacting particle systems, demonstrating that tokens, when modeled as particles, exhibit clustering under certain conditions on the Key, Query, and Value matrices. These works primarily focus on full (mean-field) attention mechanisms, where each token can interact with every other token. Building upon this foundation, our research extends the analysis to causal attention mechanisms, wherein each token is restricted to interact only with preceding tokens. This distinction is crucial, as causal attention is prevalent in Transformer models employed in generative AI and known as decoder architectures.
@@
Our contributions. Our main theoretical result establishes asymptotic clustering of tokens for causal self-attention transformer modeled as an interacting particles system on the sphere (Theorem 4.1). While mathematically accurate, this asymptotic collapse to a single cluster is seldom observed numerically. Instead, particles collapse to multiple clusters and stay in this configuration for a very long time (see Fig. 2 for a representative example) — such meta-stable states were already alluded to in Geshkovski et al., 2023b and their study was recently initiated in Koubbi et al., (2024). In Section 5 we describe such meta-stable states using analogy with the Rényi parking process (Lemma 5.1, Theorem 5.2). Additionally, Theorem 5.2 covers asymptotic clustering of tokens for causal self-attention with additional cross-attention component. Moreover, we predict that, akin to linear dynamical systems, the most important factors that qualitatively describe final particles configuration both in causal and full-attention cases are the eigenvalue of the Value matrix $V$ with the largest real part $\lambda_{\max}$ and its eigenspace $L$, while Query and Key matrices $Q,K$ and temperature parameter $\beta$ do not matter. Our conjectured atlas of possible meta-stable configurations is listed in Table 1. We prove the result stated as the first line of this table, namely that particles eventually collapse into a point when $V=I_{d}$ in Theorem 4.1. We remark that assumptions of Theorem 4.1 are much weaker than for the similar results in the full-attention case Geshkovski et al., 2023b , in particular we put no constraints on $K,Q$ or $\beta$. This work is a combination of rigorous mathematical results and non-trivial predictions based on analytical insights and numerical simulations. We summarize all limitations in Section 6.
@@
Related work. Our work builds upon the framework of Geshkovski et al., 2023b ; Geshkovski et al., 2023a where clustering properties of transformers are analyzed as systems of particles interacting on the sphere. Specifically, Geshkovski et al., 2023b proved that encoder-only (i.e. unmasked) self-attention with (post) LayerNorm leads to tokens clustering to a single point, in the limit of number of layers going to infinity. This phenomenon is also known as consensus in the related literature of multi-agent systems Markdahl et al., (2017); Criscitiello et al., (2024) and Kuramoto oscillators Strogatz, (2000); Abdalla et al., (2022). Work Geshkovski et al., 2023b in turn expands on the original perspective brought forward by Sander et al., (2022) that identify the self-attention layer as a measure-to-measure map, see also Vuckovic et al., (2021). More recently, Castin et al., (2024) studied the smoothness of this map in a framework that also covers causal attention. This work introduces a clever reparametrization that allows them to recast causal attention as mean-field dynamics, akin to their full attention counterpart. Using various approximations, Cowsik et al., (2024) were able to study a more realistic architecture that also includes MLP layers and produce accurate predictions for the final configuration of particles. This setup was further investigated by Agrachev and Letrouit, (2024) from a geometric control perspective. We note also that clustering in the absence of a residual connection (replace $\dot{x}_{k}(t)$ with $x_{k}(t+1)$ in (SA)) was established in Wu et al., (2023). Additional effects of the residual connection are studied in Dong et al., (2021) and Zhang et al., (2024).
@@ 2   Causal attention
Before describing our model of causal attention dynamics, we review the idea of Geshkovski et al., 2023b for modeling the full attention dynamics. In that work, the evolution of representations of tokens through the layers is modeled as a system of $n$ coupled Ordinary Differential Equations (ODEs) describing dynamics of a system of particles $x_{1}(t),\ldots,x_{n}(t)$. A brief part of their derivation of the dynamics from the transformers architecture is written in Section A.1. The particle position $x_{k}(t)$ corresponds to representation of the $k$-th token at layer $t$ (where for convenience, $t$ is allowed to take non-integer values) and due to RMSNorm the particles are forced to live on a unit sphere $\mathbb{S}^{d-1}$. (RMSNorm layer usually also includes a multiplication by a trainable diagonal matrix $D$, but the effect of this step can be equivalently achieved by multiplying $K,Q,V$ matrices by $D$.) These ODEs are parametrized by three matrices, known as the query $Q$, the key $K$ and the value $V$, respectively, and that are assumed to be square $d\times d$ matrices. More specifically, token $k$ evolves according to
@@
is a normalizing factor. Note that the dynamics of the $k$-th token depend on the positions of all tokens $j\in[n]$, which is a landmark characteristic of full attention leading to the so-called mean-field dynamics studied in Geshkovski et al., 2023b ; see also Geshkovski et al., 2023a ; Castin et al., (2024); Paul and Trélat, (2024).
@@ 4   Final Configuration
The system of $n$ tokens that we are studying is far more complicated than for a single token. Even establishing convergence to some point as $t\to\infty$ is challenging. In Geshkovski et al., 2023b , similar models were analyzed analytically by noticing that the dynamical system has the structure of the gradient flow of some potential function:
@@
We have already understood the evolution of the first token. In this section, we do two things. First, we describe, based on our analytical and numerical insights, conjectures about the asymptotic configuration $x(t)$ for $t\to\infty$. The surprising result here is that only the spectral properties of $V$ (and not $K$ or $Q$) affect asymptotics. Second, we rigorously prove convergence to a single point for the special case of $V=I_{d}$. We note that unlike the proof in Geshkovski et al., 2023b (see also Markdahl et al., (2017); Criscitiello et al., (2024)), our result works for all $K$ and $Q$ matrices, while the proof in Geshkovski et al., 2023b works only for $Q^{\top}K=V$ and Markdahl et al., (2017); Criscitiello et al., (2024) is restricted to $Q^{\top}K=I_{d}$.
@@ 5   Meta-stable clustering
As we discussed earlier, perhaps the most fascinating discovery of Geshkovski et al., 2023b is the existence of meta-stable clusters in the full-attention dynamics. It turns out that the same phenomenon persists in the causal-attention dynamics that we study here.
@@
From a theoretical perspective, understanding meta-stable clustering presents significant challenges, as traditional techniques for asymptotic analysis—such as those used in our Theorem 4.1—prove insufficient. Recent work on full attention transformers has made partial progress in this direction. Koubbi et al., (2024) demonstrated that when self-attention dynamics approach a nearly clustered state, they will converge to a tightly clustered configuration and remain stable for an exponential time period. Complementing this, Bruno et al., (2024) proved that tokens initialized near a uniform distribution on the sphere will spontaneously organize into a loosely clustered state. However, the bounds on the clustering tightness in this second line of work are not sufficient to trigger the convergence conditions required by Koubbi et al., (2024)’s theorem.
This Section presents a fundamental discovery regarding the identification of cluster centers in causal-attention dynamics. We establish three key claims: First, we demonstrate that initialization irregularities generate distinctive tokens, termed Rényi parking centers, which evolve into meta-stable cluster nuclei. While this phenomenon is primarily supported by numerical evidence (Fig. 3), it provides crucial insight into the clustering mechanism. Second, we prove that a subset of these special tokens, called strong Rényi centers, maintains near-stationarity over extended time periods (Lemma 5.1). Both Rényi and strong Rényi centers occur with frequency $\Theta(\beta^{\frac{d-1}{2}})$, confirming the $\sqrt{\beta}$ scaling predicted for $d=2$ by Geshkovski et al., 2023b ; see also Koubbi et al., (2024); Bruno et al., (2024). Third, we establish in Theorem 5.2 that as $t\to\infty$, all remaining tokens will converge to the vicinity of one of these stationary tokens, completing the meta-stable clustering process.
@@ A.1   From Transformers to ODEs
The derivation of the equation (SA) was thoroughly described in Geshkovski et al., 2023b , but for completeness, we briefly repeat it here to explain how the problem arises.
@@ Appendix B Final configuration
This system is autonomous, so we first explore its critical points and their stability. For autonomous systems with established convergence, it is well-known that for any absolutely continuous initialization, the limiting point is strongly unstable with probability zero (see (Shub,, 2013, Thm. III.7, Ex. III.3) and (Geshkovski et al., 2023b, , Lemma B.1)). Note that the proof in Geshkovski et al., 2023b is stated for gradient ascent dynamics but it readily extends to any smooth autonomous dynamics on a compact Riemannian manifold.
@@ Appendix C Meta-stable clustering
We know that the limiting point is critical and for almost any initialisation it is not strongly unstable. This is a well-known fact that for autonomous systems convergence to a strongly unstable point happens with probability zero, that we used in Section B.2. One could find it in (Geshkovski et al., 2023b, , Lemma B.1), based on the center manifold theorem (Shub,, 2013, Thm. III.7, Ex. III.3). Therefore, from the assumption on critical points, the limiting point is strongly stable. Then, the convergence happens exponentially fast, because locally the whole neighbourhood of a strongly stable point is its stable manifold $W^{s}_{loc}$, see (Shub,, 2013, Thm. III.7, Ex. III.3). ∎

### v2

@@ Abstract
This work presents a modification of the self-attention dynamics proposed by Geshkovski et al., (2023) to better reflect the practically relevant, causally masked attention used in transformer architectures for generative AI. This modification translates into an interacting particle system that cannot be interpreted as a mean-field gradient flow. Despite this loss of structure, we significantly strengthen the results of Geshkovski et al., (2023) in this context: While previous rigorous results focused on cases where all three matrices (Key, Query, and Value) were scaled identities, we prove asymptotic convergence to a single cluster for arbitrary key-query matrices and a value matrix equal to the identity. Additionally, we establish a connection to the classical Rényi parking problem from combinatorial geometry to make initial theoretical steps towards demonstrating the existence of meta-stable states.
@@ 1   Introduction
Following Sander et al., (2022), recent studies by Geshkovski et al., 2024b and Geshkovski et al., (2023) have proposed a mathematical framework to analyze Transformers as interacting particle systems, demonstrating that tokens, when modeled as particles, exhibit clustering under certain conditions on the Key, Query, and Value matrices. These works primarily focus on full (mean-field) attention mechanisms, where each token can interact with every other token. Building upon this foundation, our research extends the analysis to causal attention mechanisms, wherein each token is restricted to interact only with preceding tokens. This distinction is crucial, as causal attention is prevalent in Transformer models employed in generative AI and known as decoder architectures.
@@
Our contributions. Our main theoretical result establishes asymptotic clustering of tokens for causal self-attention transformer modeled as an interacting particles system on the sphere (Theorem 4.1). While mathematically accurate, this asymptotic collapse to a single cluster is seldom observed numerically. Instead, particles collapse to multiple clusters and stay in this configuration for a very long time (see Fig. 2 for a representative example) — such meta-stable states were already alluded to in Geshkovski et al., (2023) and their study was recently initiated in Geshkovski et al., 2024a . In Section 5 we describe such meta-stable states using analogy with the Rényi parking process (Lemma 5.1, Theorem 5.2). Additionally, Theorem 5.2 covers asymptotic clustering of tokens for causal self-attention with additional cross-attention component. Moreover, we predict that, akin to linear dynamical systems, the most important factors that qualitatively describe final particles configuration both in causal and full-attention cases are the eigenvalue of the Value matrix $V$ with the largest real part $\lambda_{\max}$ and its eigenspace $L$, while Query and Key matrices $Q,K$ and temperature parameter $\beta$ do not matter. Our conjectured atlas of possible meta-stable configurations is listed in Table 1. We prove the result stated as the first line of this table, namely that particles eventually collapse into a point when $V=I_{d}$ in Theorem 4.1. We remark that assumptions of Theorem 4.1 are much weaker than for the similar results in the full-attention case Geshkovski et al., (2023), in particular we put no constraints on $K,Q$ or $\beta$. This work is a combination of rigorous mathematical results and non-trivial predictions based on analytical insights and numerical simulations. We summarize all limitations in Section 6.
@@
Related work. Our work builds upon the framework of Geshkovski et al., (2023); Geshkovski et al., 2024b where clustering properties of transformers are analyzed as systems of particles interacting on the sphere. Specifically, Geshkovski et al., (2023) proved that encoder-only (i.e. unmasked) self-attention with (post) LayerNorm leads to tokens clustering to a single point, in the limit of number of layers going to infinity. This phenomenon is also known as consensus in the related literature of multi-agent systems Markdahl et al., (2017); Criscitiello et al., (2024) and Kuramoto oscillators Strogatz, (2000); Abdalla et al., (2022). Work Geshkovski et al., (2023) in turn expands on the original perspective brought forward by Sander et al., (2022) that identify the self-attention layer as a measure-to-measure map, see also Vuckovic et al., (2021). More recently, Castin et al., (2024) studied the smoothness of this map in a framework that also covers causal attention. This work introduces a clever reparametrization that allows them to recast causal attention as mean-field dynamics, akin to their full attention counterpart. Using various approximations, Cowsik et al., (2024) were able to study a more realistic architecture that also includes MLP layers and produce accurate predictions for the final configuration of particles. This setup was further investigated by Agrachev and Letrouit, (2024) from a geometric control perspective. We note also that clustering in the absence of a residual connection (replace $\dot{x}_{k}(t)$ with $x_{k}(t+1)$ in (SA)) was established in Wu et al., (2023). Additional effects of the residual connection are studied in Dong et al., (2021) and Zhang et al., (2024).
@@ 2   Causal attention
Before describing our model of causal attention dynamics, we review the idea of Geshkovski et al., (2023) for modeling the full attention dynamics. In that work, the evolution of representations of tokens through the layers is modeled as a system of $n$ coupled Ordinary Differential Equations (ODEs) describing dynamics of a system of particles $x_{1}(t),\ldots,x_{n}(t)$. A brief part of their derivation of the dynamics from the transformers architecture is written in Section A.1. The particle position $x_{k}(t)$ corresponds to representation of the $k$-th token at layer $t$ (where for convenience, $t$ is allowed to take non-integer values) and due to RMSNorm the particles are forced to live on a unit sphere $\mathbb{S}^{d-1}$. (RMSNorm layer usually also includes a multiplication by a trainable diagonal matrix $D$, but the effect of this step can be equivalently achieved by multiplying $K,Q,V$ matrices by $D$.) These ODEs are parametrized by three matrices, known as the query $Q$, the key $K$ and the value $V$, respectively, and that are assumed to be square $d\times d$ matrices. More specifically, token $k$ evolves according to
@@
is a normalizing factor. Note that the dynamics of the $k$-th token depend on the positions of all tokens $j\in[n]$, which is a landmark characteristic of full attention leading to the so-called mean-field dynamics studied in Geshkovski et al., (2023); see also Geshkovski et al., 2024b ; Castin et al., (2024); Paul and Trélat, (2024).
@@ 4   Final Configuration
The system of $n$ tokens that we are studying is far more complicated than for a single token. Even establishing convergence to some point as $t\to\infty$ is challenging. In Geshkovski et al., (2023), similar models were analyzed analytically by noticing that the dynamical system has the structure of the gradient flow of some potential function:
@@
We have already understood the evolution of the first token. In this section, we do two things. First, we describe, based on our analytical and numerical insights, conjectures about the asymptotic configuration $x(t)$ for $t\to\infty$. The surprising result here is that only the spectral properties of $V$ (and not $K$ or $Q$) affect asymptotics. Second, we rigorously prove convergence to a single point for the special case of $V=I_{d}$. We note that unlike the proof in Geshkovski et al., (2023) (see also Markdahl et al., (2017); Criscitiello et al., (2024)), our result works for all $K$ and $Q$ matrices, while the proof in Geshkovski et al., (2023) works only for $Q^{\top}K=V$ and Markdahl et al., (2017); Criscitiello et al., (2024) is restricted to $Q^{\top}K=I_{d}$.
@@ 5   Meta-stable clustering
As we discussed earlier, perhaps the most fascinating discovery of Geshkovski et al., (2023) is the existence of meta-stable clusters in the full-attention dynamics. It turns out that the same phenomenon persists in the causal-attention dynamics that we study here.
@@
From a theoretical perspective, understanding meta-stable clustering presents significant challenges, as traditional techniques for asymptotic analysis—such as those used in our Theorem 4.1—prove insufficient. Recent work on full attention transformers has made partial progress in this direction. Geshkovski et al., 2024a demonstrated that when self-attention dynamics approach a nearly clustered state, they will converge to a tightly clustered configuration and remain stable for an exponential time period. Complementing this, Bruno et al., (2024) proved that tokens initialized near a uniform distribution on the sphere will spontaneously organize into a loosely clustered state. However, the bounds on the clustering tightness in this second line of work are not sufficient to trigger the convergence conditions required by Geshkovski et al., 2024a ’s theorem.
This Section presents a fundamental discovery regarding the identification of cluster centers in causal-attention dynamics. We establish three key claims: First, we demonstrate that initialization irregularities generate distinctive tokens, termed Rényi parking centers, which evolve into meta-stable cluster nuclei. While this phenomenon is primarily supported by numerical evidence (Fig. 3), it provides crucial insight into the clustering mechanism. Second, we prove that a subset of these special tokens, called strong Rényi centers, maintains near-stationarity over extended time periods (Lemma 5.1). Both Rényi and strong Rényi centers occur with frequency $\Theta(\beta^{\frac{d-1}{2}})$, confirming the $\sqrt{\beta}$ scaling predicted for $d=2$ by Geshkovski et al., (2023); see also Geshkovski et al., 2024a ; Bruno et al., (2024). Third, we establish in Theorem 5.2 that as $t\to\infty$, all remaining tokens will converge to the vicinity of one of these stationary tokens, completing the meta-stable clustering process.
@@ A.1   From Transformers to ODEs
The derivation of the equation (SA) was thoroughly described in Geshkovski et al., (2023), but for completeness, we briefly repeat it here to explain how the problem arises.
@@ Appendix B Final configuration
This system is autonomous, so we first explore its critical points and their stability. For autonomous systems with established convergence, it is well-known that for any absolutely continuous initialization, the limiting point is strongly unstable with probability zero (see (Shub,, 2013, Thm. III.7, Ex. III.3) and (Geshkovski et al.,, 2023, Lemma B.1)). Note that the proof in Geshkovski et al., (2023) is stated for gradient ascent dynamics but it readily extends to any smooth autonomous dynamics on a compact Riemannian manifold.
@@ Appendix C Meta-stable clustering
We know that the limiting point is critical and for almost any initialisation it is not strongly unstable. This is a well-known fact that for autonomous systems convergence to a strongly unstable point happens with probability zero, that we used in Section B.2. One could find it in (Geshkovski et al.,, 2023, Lemma B.1), based on the center manifold theorem (Shub,, 2013, Thm. III.7, Ex. III.3). Therefore, from the assumption on critical points, the limiting point is strongly stable. Then, the convergence happens exponentially fast, because locally the whole neighbourhood of a strongly stable point is its stable manifold $W^{s}_{loc}$, see (Shub,, 2013, Thm. III.7, Ex. III.3). ∎